# Meritocracy and Its Discontents: Long-run Effects of Repeated School Admission Reforms

### v1

@@ Abstract
What happens if selective colleges change their admission policies? We answer this question by analyzing the world’s first introduction of nationally centralized meritocratic admissions in the early twentieth century. We find a persistent meritocracy-equity tradeoff. Compared to the decentralized system, the centralized system admitted more high-achieving applicants, producing a greater number of top elite bureaucrats in the long run. However, this impact came at the distributional cost of equal regional access to elite higher education and career advancement. Several decades later, the meritocratic centralization increased the number of urban-born career elites (e.g., top income earners) relative to rural-born ones.
@@ 1 Introduction
College and school admission processes vary across time and place. American college admissions use a decentralized system where each college makes its own admissions often based on opaque evaluation criteria. In contrast, China’s public college admissions illustrate a regionally-integrated, single-application, and single-offer system using a transparent admission criterion. How do different admissions institutions affect students’ education and future career? What are the impacts on the national production of high-skilled individuals and their distributions?
@@
In this paper, we study the impacts of using centralized and meritocratic admissions. We do so by combining a series of natural experiments in history, newly assembled historical data, and theoretical insights. Our investigations reveal the pros and cons of centralized meritocratic admissions, especially a tradeoff between meritocracy and equal access to selective higher education and career achievements.
@@
Our empirical setting is the first known transition from decentralized to nationally-centralized school admissions. At the end of the 19th century, to modernize its higher education system, the Japanese government set up National Higher Schools (roughly equivalent to today’s colleges) as an exclusive entry point to the most prestigious tertiary education. [^1] These schools were transformed into parts of today’s national universities after WWII. For example, the National Higher School in Tokyo became the first two years of the University of Tokyo. These schools later produced many of the most influential members of the society, including several Prime Ministers, Nobel Laureates, and founders of global companies like Toyota. Their selective and prestigious nature and social influences had similarities to Oxbridge in the UK, Grande Ecoles in France, and the Ivy League in the US. [^2] We study top elites on both substantive and pragmatic grounds. The individual holdings of prestigious economic, political, and cultural positions are skewed and concentrated on a tiny minority of the population who have big influences on the economy and society. For their social importance, we also have relatively rich historical data that allows us to trace back their careers from humble beginnings as students to wide societal recognition and acclaim.
@@
Acceptance into these schools was merit-based, using annual entrance examinations. Initially, the government let each school run its own exam and admissions based on exam scores, similar to many of today’s decentralized K-12 and college admissions. Under this decentralized system, the schools typically held exams on the same day so that each applicant could apply for only one school. [^3] Similar restrictions on the number of applications exist today in the college admission systems of Italy, Nigeria, and the UK. As a result, many high achieving applicants applied to too competitive schools and failed to enter any school.
@@
This problem motivated the government to introduce a centralized system in order to allocate the limited school seats to the highest-achieving students. In the new system introduced at the turn of the 20th century, applicants were allowed to rank multiple schools in the order of their preferences and take a single unified exam. [^4] As shown later, the defining feature of the centralized system was to allow applicants to list multiple schools, not the use of a single unified exam. Given their preferences and exam scores, each applicant is assigned to a school (or none if unsuccessful) based on a computational algorithm. The algorithm was a mix of the Deferred Acceptance and Immediate Acceptance (Boston) algorithms combined with a meritocracy principle that only assigns the highest-scoring applicants to any school. To our knowledge, this instance is the first nationwide use of any matching algorithm. [^5] The earliest known large-scale use of the Boston algorithm is the assignment of medical residents to hospitals in New York City in the 1920s (Roth, 1990). The oldest known national use of the Deferred Acceptance algorithm is the National Resident Matching Program (NRMP) in the 1950s (Roth, 1984). See Abdulkadiroğlu and Sönmez (2003) for the details of these algorithms in school admission contexts. Furthermore, for reasons detailed below, the government later re-decentralized and re-centralized the system several times, producing multiple natural experiments for studying the consequences of the different systems. [^6] During 1900–1930, there were three periods of centralization (1902–07, 1917–18, and 1926–27). We exploit these bidirectional institutional changes to identify the short- and long-run impacts of the meritocratic centralization.
A stylized theoretical model predicts the short-run effects of the reforms on application behavior and admission outcomes. Consistent with its stated goal, the model confirms that the centralized system produces more meritocratic school seat allocations than the decentralized system. It also predicts that centralization would cause applicants to apply to more selective schools and make more inter-regional applications.
To test these predictions and analyze the effects of meritocratic centralization, we newly collect and digitize data about application and enrollment. The data is based on administrative school records, Government Gazettes, and Ministry of Education Yearbooks. Our data records each student’s regional origin, which is a strong predictor of socioeconomic status both in our empirical setting and other settings like the contemporary US.
@@
We find that meritocratic centralization had large and immediate effects on application and enrollment. First, consistent with the theoretical predictions, centralization caused stark strategic responses in application behavior. In particular, strategic incentives in the centralized system led applicants across the country to more often rank the most selective school first. This effect and the option to apply to multiple schools made students enroll at schools further away from their regional origin. The reform thus made the elite higher education market more meritocratic and regionally integrated.
@@
The gains from centralization were not equally distributed, however. Because high-achieving students were disproportionately located in urban areas (mainly Tokyo), the centralized system caused a greater number of urban applicants to be admitted to schools in rural areas, typically after being rejected by their first choice schools. [^7] We use the nomenclature of “urban” and “rural” schools, but note that “rural” schools were located in regional cities rather than in the countryside. As a result, urban high-achievers crowded out rural applicants; the number of entrants to any national elite school coming from the urban area increased by about 10% during centralization. [^8] It is also empirically true that the centralized system made a greater number of rural applicants apply to and enter urban schools. The centralized system thus increased regional mobility across the country. But their net effects are such that urban high-achievers crowded out rural applicants.
Historical documents suggest that this distributional consequence upset rural schools and communities. Partly as a result of such rural discontents, the government went back and forth between decentralized and centralized systems, finally settling on a decentralized scheme. This series of bidirectional reforms enables us to identify the causal effects of meritocratic centralization more precisely than a usual, single policy change would. [^9] We provide evidence that the reform timings appear exogenous in that they are not correlated with other social and institutional outcomes.
Our main results are about the long-run effects of the centralized system. We start by looking at the persistence of the distributional impacts. The short-run analysis finds that centralization let urban areas disproportionately gain school access relative to rural areas. This result motivates a difference-in-differences analysis that compares long-term career outcomes of urban vs rural-born individuals by each cohort’s exposure to the centralized system. Our long-run career data comes from the Personnel Inquiry Records (PIR) published in 1939, more than thirty years after the first episode of meritocratic centralization in 1902–07. This data provides a list of distinguished individuals (e.g., high-income earners, national medal recipients, high-ranking politicians, corporate executives) along with their biographical information. We provide extensive investigations about the quality of the data. [^10] We find that the data covers a large fraction of the national population of elites (e.g., 53% of the top 0.01% income earners) by comparing our data with national administrative statistics. We also find no systematic variation in the sampling rates across prefectures, consistent with our assumption that sample selection bias is uncorrelated with the prefecture-cohort variation we use.
@@
Distributional effects of meritocratic centralization turn out to be persistent. Almost four decades later, relative to the decentralized system, the centralized system produced a greater number of top income earners, prestigious medal recipients, and other elite professionals who came from urban areas compared to rural areas. The number of urban-born career elites increased by 10–20% for the cohorts exposed to meritocratic centralization. We also obtain suggestive evidence that the centralized system increased the number of career elites living in urban areas as adults in the long run.
@@
We conclude the long-run analysis by examining the overall productive efficiency implication of meritocratic centralization for the whole country. We focus on a specific subgroup of career elites — top bureaucrats. Top bureaucrats were an exemplar category of career elites in our empirical context, as the best and brightest graduates from the top colleges competed to become bureaucrats. Using the comprehensive administrative data of all individuals who passed the national exams to become higher civil officials, we compare the national number of bureaucrats promoted to top ranks between cohorts exposed to the centralized admissions and cohorts exposed to the decentralized admissions. We find that the number of top-ranking bureaucrats in the centralized cohorts was 15% greater than that in the decentralized cohorts, suggesting that centralized cohorts won promotion competitions over close decentralized cohorts. Meritocratic centralization thus produced a larger number of high-quality bureaucratic elites for the whole country.
@@
Overall, our findings reveal an equity-meritocracy tradeoff both in the short and long run. On the one hand, the meritocratic centralization of school admissions achieved the goal of rewarding applicants with higher academic performance in the short run. This meritocracy resulted in a greater number of top bureaucratic elites produced by the exposed cohorts in the long run. On the other hand, this meritocracy came at the equity cost of urban-born high-achieving applicants overwhelming rural-born ones. The design of admission rules therefore affects the production and distribution of highly educated and skilled individuals, which is an important determinant of economic growth and inequality (Glaeser, 2011; Moretti, 2012).
@@
Literature. Our analysis contributes to the empirical literature that uses policy reforms to study the impacts of the design of admission systems (Abdulkadiroğlu et al., 2006; Abdulkadiroğlu, Agarwal and Pathak, 2017; Chen, Jiang and Kesten, 2020; Machado and Szerman, 2021; Grenet, He and Kübler, 2022; Kapor, Karnani and Neilson, 2023). [^11] Other studies measure the effects of selective schools conditional on a particular admission system (Dale and Krueger, 2002; Hastings, Neilson and Zimmerman, 2013; Pop-Eleches and Urquiola, 2013; Deming et al., 2014; Kirkeboen, Leuven and Mogstad, 2016; Clark and Del Bono, 2016; Abdulkadroǧlu et al., 2017; Zimmerman, 2019; Abdulkadiroglu et al., 2021; Michelman, Price and Zimmerman, 2021; Chetty, Deming and Friedman, 2023). While these prior studies focus on the short-run effects (on application, enrollment, academic achievement, and welfare), we estimate the long-run effects by taking advantage of bidirectional, repeated policy changes in history. [^12] With its interest in long-run effects, this paper also relates to studies of the long-term effects of educational reforms (e.g., Duflo, 2001; Meghir and Palme, 2005). These studies focus on the effects of expanding resources (such as school constructions and compulsory education extensions), while we investigate the effects of changing resource allocation mechanisms given a fixed amount of resources. This use of bidirectional policy changes echoes other studies with similar identification strategies (Niederle and Roth, 2003). Our analysis of the meritocracy vs equity tension has some similarities to ongoing policy discussions on affirmative actions, which are surveyed in Arcidiacono and Lovenheim (2016). Recent contributions in this area include Bleemer (2020, 2022), Kapor (2020) and Otero, Barahona and Dobbin (2021). [^13] Kamada and Kojima (2015) and Agarwal (2017) study regional equity in other matching markets.
@@
From a broader historical perspective, this paper relates to the literature that investigates the long-term effects of institutions (Nunn, 2009), particularly resource allocation mechanisms (Bleakley and Ferrie, 2014, 2016). Our analysis is also related to Bai and Jia (2016), who examine political consequences of the abolition of a meritocratic elite recruitment system (civil service exam) in early twentieth-century China.
@@
The next section provides the institutional background. Section 3 examines the short-term effects of meritocratic centralization, while Section 4 analyzes the long-term impacts. Finally, Section 5 summarizes our findings and outlines future directions.
@@ 2 Background
One major trend in modern college admissions is a growing degree of centralization with transparent admission criteria. Today, over 30 countries use regionally- or nationally-integrated, single-application and single-offer college admissions. Centralized college admissions are used in all continents except North America (Appendix Figure A.1). These systems have well-specified admission criteria, mixing meritocratic achievement elements (such as GPA and entrance exams), and affirmative action and other priority considerations.
Before the turn of the 20th century, however, no country used such a centralized system. Even today, many countries, including the US, continue to use decentralized systems. Such decentralized schemes tend to come with less transparent criteria for ranking applicants, as illustrated by recent court cases against American universities. Similar observations also apply to K-12 school admissions. How does the centralization of college and school admissions affect students’ application behavior, enrollment outcomes, and future careers? Understanding the costs and benefits of meritocratic centralization is the goal of our paper.
@@ 2.1 Bidirectional Admission Reforms
To evaluate the impacts of centralized admission systems, we take advantage of unique historical episodes in early twentieth-century Japan. Since the opening of the country in the mid 19th century, to catch up with Western knowledge, science, and technologies, education reforms became a central part of modernization efforts by the Japanese government. In 1894, the government set up a national higher education system consisting of one Imperial University and five National Higher Schools. By 1908, the system was expanded to four Imperial Universities (Tokyo, Kyoto, Tohoku, and Kyushu) and eight National Higher Schools located across different areas in Japan (First in Tokyo, Second in Sendai, Third in Kyoto, Fourth in Kanazawa, Fifth in Kumamoto, Sixth in Okayama, Seventh in Kagoshima, and Eighth in Nagoya, named after the order of establishment; see Appendix Figure A.2 for their locations). We refer to these eight National Higher Schools as Schools 1–8 for short. [^14] The number of higher education institutions increased after 1918, as the government permitted not only national but also local public and private higher schools and universities. See Appendix Section A.1 and Table A.1 for details. In our empirical analyses, we control for the number of National Higher Schools as well as other characteristics of higher education institutions. We focus on Schools 1–8.
@@
Schools 1–8 served as an exclusive entry point to Imperial Universities, the most prestigious tertiary education. Virtually all graduates of Schools 1–8 were admitted to these universities without further selection well into the 1920s. Imperial University graduates were also partially or wholly exempted from the Higher Civil Service Examinations and other selective national qualification exams to become high-ranking administrators, diplomats, judges, and physicians. As a result, entering Schools 1–8 was considered as a passport into the elite class. Indeed, Schools 1–8 produced highly distinguished and influential individuals, including several Prime Ministers, Nobel Laureates, world-leading mathematicians, renowned novelists, and founders of global companies like Toyota. To apply to these schools, one must be male aged 17 or older and have completed a five-year middle school. [^15] The eligibility was changed in 1919 to males aged 16 or older who have completed the fourth year of middle school. Schools 1–8 admitted fewer than 2,300 students each year throughout 1900–1930, constituting less than 0.5% of the cohort of males aged 17. More descriptive statistics are presented in Appendix Table A.2.
@@
Among Schools 1-8, School 1 in Tokyo was considered the most prestigious due to its location in the capital and its geographical proximity to Tokyo Imperial University (today’s University of Tokyo). The next most prestigious was School 3 in Kyoto, the ancient capital of Japan. By contrast, located in a remote southwest region, Schools 5 and 7 were considered the least prestigious among all schools. Consequently, the schools differed substantially in their popularity and selectiveness. [^16] Tuition was mostly uniform across Schools 1-8.
@@
The admission to Schools 1–8 was merit-based and determined by annual entrance exams. Initially, the government let each school administer its own exam and admissions. Schools 1–8 typically held their exams on the same day so that each applicant could only apply to one school. Following the convention in the literature (Che and Koh, 2016; Hafalir et al., 2018), we call this system “decentralized admissions” or Dapp for short. The single choice aspect of Dapp captures an essential feature of decentralization, which incentivizes each applicant to self-select into an appropriate school by comparing the selectivity of schools with his own standing.
@@
Under Dapp, however, many high-achieving students were rejected by popular schools, while lower-achieving students entered less popular schools. For the government, who wanted to send the most promising students to selective higher education, the decentralized system appeared inefficient. [^17] Applicants who failed to enter Schools 1-8 chose either (1) to retake the exam in the following year, (2) to enter local public or private higher educational institutions, or (3) to give up receiving higher education. The Education Minister criticized the decentralized system as follows:
@@
[Under the decentralized system] many applicants are rejected by Schools 1 and 3 [in Tokyo and Kyoto], which attract a large number of high ability applicants, despite the fact that these applicants have superior academic performance to that of applicants admitted to other rural schools. (…) As a result, hundreds of applicants with sufficiently high academic ability to enter rural schools are idly wasting another year [to retake the exam]. This is not only a pity for them, but also a loss for the country.” (Educational Review No.1146, p.21, published in February 15, 1917)
@@
To solve this problem, in 1901, the government first asked all schools to unify their entrance exams to a single one, while maintaining decentralized admission decisions. In 1902, the government launched a centralized system in which applicants apply for multiple schools, submit their preferences over schools, and take the unified exam. Based on their exam scores and preferences, applicants were then assigned to a school (or no school if unsuccessful) by an assignment algorithm announced ex ante. Each year, the government announced application procedures in April, three months before the exam in July, as a public notice in Government Gazette. We call this system “centralized admissions” or Capp for short.
@@
The centralized assignment algorithm operates as follows. [^18] This is the assignment algorithm used in 1917. See Appendix Figure A.3 for a reprint of the original assignment algorithm published in April 1917. Strictly speaking, the centralized assignment algorithm was introduced first in 1902, slightly modified in 1903 and 1917, and revised further in 1926.
@@
    For applicants selected in (1) and for whom the school of his second choice is already filled at the end of (3), then assign each applicant to the school of his third choice or below, repeating the same procedure as (3).
@@
Written more than a century ago in natural language, the rule description is mathematically precise. The above algorithm imposes meritocracy up front in which only top-scoring applicants were considered for admission regardless of their preferences (Step (1)). Capp is meritocratic in the sense that the test score distribution among assigned students under Capp first order stochastically dominates that under any other mechanisms, including Dapp. This step selects only applicants who would be admitted by any school under the Deferred Acceptance algorithm (Abdulkadiroğlu and Sönmez, 2003), one of the most widely used algorithms in today’s college and selective K-12 admissions (Appendix B). These applicants are then assigned to one of Schools 1–8 using the Immediate Acceptance (Boston) algorithm (Steps (2) to (4)). This algorithm is therefore a variant of the Immediate Acceptance algorithm with a meritocracy constraint, making it closer to the Deferred Acceptance algorithm. To the best of our knowledge, this is the world’s first recorded nation-wide use of any assignment algorithm.

### v2

@@ Abstract
What happens if selective colleges change their admission policies? We study this question by analyzing the world’s first implementation of nationally centralized meritocratic admissions in the early twentieth century. We find a persistent meritocracy-equity tradeoff. Compared to the decentralized system, the centralized system admitted more high-achievers and produced more occupational elites (such as top income earners) decades later in the labor market. This gain came at a distributional cost, however. Meritocratic centralization also increased the number of urban-born elites relative to rural-born ones, undermining equal access to higher education and career advancement.
@@ 1 Introduction
One major trend in college admissions around the world is a growing degree of centralization. Today, over thirty countries use regionally- or nationally-integrated, single-application and single-offer college admissions (Appendix Figure A.1). These systems have well-specified admission criteria, mixing meritocratic achievement elements (such as GPA and entrance exams) and other priority considerations. Before the turn of the 20th century, however, no country used such a centralized system. How does the centralization of admission systems affect students’ life trajectories? What are their impacts on the national production of highly skilled individuals and their composition? A key challenge in studying these questions is the lack of clear policy changes and data about students’ long-run outcomes.
@@
In this paper, we study the impacts of centralized and meritocratic college admissions. Our investigations reveal their pros and cons, especially a tension between meritocracy and equal access to higher education and career achievements in the long run. We reach these findings by combining a series of natural experiments in history and newly assembled historical data that trace students over decades.
@@
Our empirical setting is the world’s first known transition from decentralized to nationally-centralized school admissions. At the end of the 19th century, to modernize its higher education system, the Japanese government set up National Higher Schools (roughly equivalent to today’s colleges) as an exclusive entry point to the most prestigious tertiary education. [^1] After WWII, National Higher Schools were transformed into parts of today’s national universities. For example, the First National Higher School in Tokyo became the first two years of the University of Tokyo. We focus on these selective colleges, as they later produced many of the most influential leaders of the society, including several Prime Ministers, Nobel Laureates, and founders of global companies like Toyota. Their prestige and social influences were similar to Oxbridge in the UK, Grandes Écoles in France, and the Ivy League in the US.
@@
Acceptance into these schools was based on annual entrance exams. Initially, the government let each school run its own exam and admissions based on exam scores, similar to many of today’s decentralized K-12 and college admissions. Under this decentralized system, the schools typically held exams on the same date so that each applicant could apply for only one school. [^2] Similar restrictions on the number of applications exist today in the college admission systems of Italy, Nigeria, and the UK. As a result, many high achieving applicants applied to and were rejected by the most competitive school, failing to enter any school.
@@
This problem motivated the government to introduce a centralized system in 1902. The new system prioritized the admission of the highest-scoring students. Specifically, applicants were asked to submit preference rankings over multiple schools and to take a single unified exam. [^3] As shown later, the defining feature of the centralized system was to allow applicants to list multiple schools, not the use of a single unified exam. Given their preference rankings and exam scores, each applicant was assigned to a school (or none if unsuccessful) based on a computational algorithm. The algorithm was a mix of the Deferred Acceptance and Immediate Acceptance (Boston) algorithms combined with a meritocracy principle that assigns only the highest-scoring applicants to any school. To our knowledge, this instance is the first nationwide use of any matching algorithm. [^4] The earliest known large-scale use of the Boston algorithm is the assignment of medical residents to hospitals in New York City in the 1920s (Roth, 1990). The oldest known national use of the Deferred Acceptance algorithm is the National Resident Matching Program (NRMP) in the 1950s (Roth, 1984). See Abdulkadiroğlu and Sönmez (2003) for the details of these algorithms in school admission contexts.
Furthermore, for reasons detailed below, the government later re-decentralized and re-centralized the system several times, producing multiple natural experiments for studying the consequences of the different systems. [^5] During 1900–1930, there were three periods of centralization in 1902–1907, 1917–1918, and 1926–1927. We exploit these bidirectional institutional changes to identify the short- and long-run impacts of meritocratic centralization. For the short-run analysis, we newly collect and digitize application and enrollment data from administrative school records and government documents, including Government Gazettes and Ministry of Education Yearbooks.
@@
We first find that meritocratic centralization had large effects on application behavior and enrollment outcomes. In particular, centralization caused applicants in all areas to be more risk-taking and more often rank the most selective school first. Applicants also made longer-distance applications, leading students to be enrolled at schools further away from their regional origin. The reform thus made the selective higher education market more meritocratic, competitive, and regionally integrated.
@@
The effect of centralization was heterogeneous, however. Because high-achieving students were disproportionately located in urban areas (mainly Tokyo), the centralized system caused a greater number of urban applicants to be admitted to schools in rural areas, typically after being rejected by their first-choice schools. [^6] We use “urban areas” to refer to a region surrounding Tokyo (as defined in Section 3). As shown later, these areas were characterized by a greater population, higher income, and better educational infrastructure. As a result, urban high-achievers crowded out rural applicants: the number of urban-born entrants to any national higher school increased by about 10% during centralization. [^7] It is also empirically true that the centralized system made a greater number of rural applicants apply to and enter urban schools. The centralized system thus increased regional mobility across the country. But their net effects are such that urban high-achievers crowded out rural applicants. Historical documents suggest that this distributional consequence upset rural schools and their local communities. Such rural discontents were a reason why the government oscillated between the centralized and decentralized systems.
Our main results are the long-run effects of the centralized system on the national production and regional distribution of occupational elites. To study the distributional consequence, we compare urban vs rural-born individuals’ long-term career outcomes by each cohort’s exposure to the centralized admission system. The career outcome data is based on the Personnel Inquiry Records (PIR) published in 1939 (more than thirty years after the first period of meritocratic centralization). The data provides a list of socially distinguished individuals—encompassing economic, political, and cultural elites—and their biographical information. [^8] We provide extensive investigations about the quality of the data (such as the coverage and sampling bias) in Section 4.1.
@@
Distributional effects of meritocratic centralization turn out to be persistent. Almost four decades after the reform, relative to the decentralized system, the centralized system produced a greater number of occupational elites (such as top income earners and top politicians and bureaucrats) who came from urban areas compared to rural areas. The number of urban-born elites increased by 10–20% for the cohorts exposed to meritocratic centralization. We also provide suggestive evidence that the centralized system increased the number of elites living in urban areas in their late adulthood. Meritocratic centralization thus impacted both the origins and destinations of highly skilled individuals.
@@
We conclude our analysis by examining the impact of meritocratic centralization on the national production of occupational elites. We first focus on a specific subgroup, i.e., top bureaucrats. Top bureaucrats were an important category of elites in our empirical context, where higher civil service was one of the most prestigious and coveted jobs due to their political and legislative influence in building the state (Shimizu, 2019). Using comprehensive administrative data of all individuals who have passed the national exams to become higher civil officials, we compare the total number of bureaucrats promoted to top ranks between cohorts exposed to the centralized admissions and cohorts exposed to the decentralized admissions. We find that the number of top-ranking bureaucrats was 15% greater in the centralized cohorts than in the decentralized cohorts, suggesting that centralized cohorts won promotion competitions over decentralized cohorts. Moreover, we find that meritocratic centralization produced a greater total number of all occupational elites listed in the PIR data, including top-income earners. Meritocratic centralization thus produced a greater number of occupational elites for the whole country.
@@
Overall, our findings show that the design of admission rules affects the production and distribution of highly educated and skilled individuals, which is an important determinant of economic growth and inequality (Glaeser, 2011; Moretti, 2012). On one hand, meritocratic centralization achieved the goal of recruiting applicants with higher academic performance. Moreover, this meritocracy produced a greater number of occupational elites in the long run. The policymaker may therefore embrace meritocratic centralization on the ground that it produces more high-quality leaders for the society.
On the other hand, this gain came at the equity cost of urban-born high-achievers overwhelming rural-born ones both in the short and long run. Meritocratic centralization expanded the urban-rural gap in the ability to produce highly skilled individuals. Meritocracy may thus dampen the society’s capability to diversify the backgrounds of individuals in leadership positions and represent a broad range of socioeconomic groups.
@@
Literature. Our analysis contributes to the empirical literature that uses policy reforms to study the impacts of the design of admission systems (Abdulkadiroğlu et al., 2006; Abdulkadiroğlu, Agarwal and Pathak, 2017; Chen, Jiang and Kesten, 2020; Machado and Szerman, 2021; Terrier, Pathak and Ren, 2021; Kapor, Karnani and Neilson, 2023; Campos and Kearns, 2024). [^9] Other studies measure the effects of selective schools conditional on particular admission systems (Dale and Krueger, 2002; Hastings, Neilson and Zimmerman, 2013; Pop-Eleches and Urquiola, 2013; Deming et al., 2014; Kirkeboen, Leuven and Mogstad, 2016; Clark and Del Bono, 2016; Abdulkadroǧlu et al., 2017; Zimmerman, 2019; Abdulkadiroglu et al., 2021; Narita, 2021; Michelman, Price and Zimmerman, 2021; Chetty, Deming and Friedman, 2023). While these prior studies focus on the short-run effects (on application, enrollment, academic achievement, and welfare), we estimate the long-run effects by taking advantage of bidirectional, repeated policy changes in history. [^10] With its interest in long-run effects, this paper also relates to studies of the long-term effects of educational reforms (e.g., Duflo, 2001; Meghir and Palme, 2005; Ichimura et al., 2024). These studies focus on the effects of expanding resources (such as school constructions and compulsory education extensions), while we investigate the effects of changing resource allocation mechanisms given a fixed amount of resources. This use of bidirectional policy changes echoes other studies with similar identification strategies (Niederle and Roth, 2003). Our analysis of the meritocracy vs equity tension has some similarities to ongoing policy discussions on affirmative actions, which are surveyed in Arcidiacono and Lovenheim (2016). Recent important contributions in this area include Bleemer (2020, 2022), Kapor (2020) and Otero, Barahona and Dobbin (2021). [^11] Kamada and Kojima (2015) and Agarwal (2017) study regional equity in other matching markets.
@@
From a broader historical perspective, this study relates to the literature that investigates the long-term effects of institutions (Nunn, 2009), particularly the effects of resource allocation mechanisms (Bleakley and Ferrie, 2014). Our analysis is also related to Bai and Jia (2016), who examine political consequences of the abolition of a meritocratic elite recruitment system (civil service exam) in early twentieth-century China.
@@
The next section provides the historical and institutional background. Section 3 examines the short-term effects of meritocratic centralization, while Section 4 analyzes the long-term impacts. Section 5 summarizes our findings and outlines future directions.
@@ 2.1 Bidirectional Admission Reforms
To analyze the impact of centralized admissions, we take advantage of unique historical episodes in early twentieth-century Japan. Since the opening of the country in the mid 19th century, to catch up with Western knowledge, science, and technologies, education reforms became a central part of modernization efforts by the Japanese government. In 1894, the government set up a national higher education system consisting of one Imperial University (four-year program) and five National Higher Schools (five-year program). By 1908, the system was expanded to four Imperial Universities (Tokyo, Kyoto, Tohoku, and Kyushu) and eight National Higher Schools located across the country. [^12] First in Tokyo, Second in Sendai, Third in Kyoto, Fourth in Kanazawa, Fifth in Kumamoto, Sixth in Okayama, Seventh in Kagoshima, and Eighth in Nagoya, named after the order of establishment; see Appendix Figure A.2 for their locations. We refer to these eight National Higher Schools as Schools 1–8 for short. [^13] The number of higher education institutions increased after 1918, as the government permitted not only national but also local public and private higher schools and universities. See Appendix Section A.1 and Table A.1 for details. In our empirical analyses, we control for the number of National Higher Schools as well as other characteristics of higher education institutions. Our analysis focuses on Schools 1–8.
@@
Schools 1–8 served as an exclusive entry point to Imperial Universities, the most prestigious tertiary education. Virtually all graduates of Schools 1–8 were admitted to these universities without further selection well into the 1920s. Imperial University graduates were also partially or wholly exempted from the Higher Civil Service Examinations and other selective national qualification exams to become high-ranking administrators, diplomats, judges, and physicians. As a result, entering Schools 1–8 was considered as a passport into the elite class. Indeed, Schools 1–8 produced highly distinguished and influential individuals, including Prime Ministers, Nobel Laureates, world-leading mathematicians, renowned novelists, and founders of global companies like Toyota. To apply to these schools, one must be male aged 17 or older and have completed a five-year middle school. [^14] The eligibility was changed in 1919 to males aged 16 or older who have completed the fourth year of middle school. Schools 1–8 admitted fewer than 2,300 students each year throughout 1900–1930, constituting less than 0.5% of the cohort of males aged 17. Descriptive statistics are presented in Appendix Table A.2.
@@
Among Schools 1–8, School 1 in Tokyo was considered the most prestigious due to its location in the capital and its geographical proximity to Tokyo Imperial University (today’s University of Tokyo). The next most prestigious was School 3 in Kyoto, the ancient capital of Japan. By contrast, located in a remote southwest region, Schools 5 and 7 were considered the least prestigious among all schools. Consequently, the schools differed substantially in their popularity and selectiveness. [^15] Tuition was mostly uniform across Schools 1–8.
@@
The admission to Schools 1–8 was based on annual entrance exams. Initially, the government let each school administer its own exam and admissions. Typically, Schools 1–8 coordinated to hold their exams on the same date so that each applicant could only apply to one school. We call this system “decentralized admissions” (or “decentralization” for short). The single choice aspect captures an essential feature of decentralization, which incentivizes each applicant to self-select into an appropriate school by comparing the selectivity of schools with his own standing (Che and Koh, 2016).
@@
Under decentralization, however, many high-achieving students were rejected by popular schools, while lower-achieving students entered less popular schools. For the government, who wanted to send the most promising students to selective higher education, the decentralized system appeared inefficient. [^16] Applicants who failed to enter Schools 1–8 chose either (1) to retake the exam in the following year, (2) to enter local public or private higher educational institutions, or (3) to give up receiving higher education. The Education Minister criticized the decentralized system as follows:
@@
[Under the decentralized system] many applicants are rejected by Schools 1 and 3 [in Tokyo and Kyoto], which attract a large number of high ability applicants, despite the fact that these applicants have superior academic performance to that of applicants admitted to other more rural schools. (…) As a result, hundreds of applicants with sufficiently high academic ability to enter rural schools are idly wasting another year [to retake the exam]. This is not only a pity for them, but also a loss for the country. [^17] A quote from Educational Review No.1146, p.21, published in February 15, 1917, right before the second introduction of the centralized system.
@@
To solve this problem, in 1901, the government first asked all schools to unify their entrance exams to a single one, while maintaining decentralized admission decisions. In 1902, the government launched a centralized system. Under the new system, applicants were allowed to apply for multiple schools by submitting their preference rankings over schools before taking the unified exam. Based on their exam scores and school preferences, applicants were then assigned to one school (or no school if unsuccessful) by an assignment algorithm announced ex ante. Each year, the government announced application procedures in April, three months before the exam in July, as a public notice in the Government Gazette. We call this system centralized admissions (or centralization for short).
@@
The centralized assignment algorithm was specified as follows. [^18] This is the assignment algorithm in 1917 (see Appendix Figure A.3 for a reprint of the original assignment algorithm published in the Government Gazette). The centralized assignment algorithm was first introduced in 1902 and reintroduced in 1917 with a slight modification.
@@
    For applicants selected in (1) and for whom the school of his second choice is already filled at the end of (3), assign each applicant to the school of his third choice or below, repeating the same procedure as (3).
@@
Written more than a century ago in natural language, the rule description is mathematically precise. The above algorithm imposes meritocracy up front in which only top-scoring applicants were considered for admission regardless of their preferences (Step (1)). The centralized system is meritocratic in the sense that the test score distribution among the assigned students first order stochastically dominates that under any other mechanisms, including the original decentralized system. Step (1) also selects the same applicants as those who would be admitted by any school under the Deferred Acceptance algorithm, one of the most widely used algorithms in today’s college and selective K-12 admissions (see Appendix B for details). These applicants are then assigned to one of Schools 1–8 using the Immediate Acceptance (Boston) algorithm (Steps (2) to (4)). This algorithm is therefore a variant of the Immediate Acceptance algorithm with a meritocracy constraint, making it closer to the Deferred Acceptance algorithm. To the best of our knowledge, this is the world’s first recorded nation-wide use of any assignment algorithm.
2.2 Political Economy of the Admission Reforms